OpenStudy (anonymous):

Find x when 2^x=16^y, 16^x=8^(y-1) Okay when I exponentiate for the first part I get y=x/3 and for the second part I get y=4x-1/3. However, I don't understand how to get one x. I don't know what the next step is. Please help me.

OpenStudy (anonymous):

\[2^x=16^y\] \[2^x=2^{4y}\] \[x=4y\]

OpenStudy (anonymous):

then \[16^x=8^{y-1}\] \[2^{4x}=2^{3(y-1)}\] \[4x=3y-3\]

OpenStudy (anonymous):

should be good from there yes?

OpenStudy (anonymous):

okay I see how you got that but in order to find x do I have to set them equal to each other? I am given answers and it is not in that form. The answers are like in fraction form.

OpenStudy (anonymous):

There is only one answer and I am confused because I am given two equations for one problem

OpenStudy (anonymous):

I have a question @khuangg , howcome on 16^x=8^(y-1) you distributed the exponent 3 in to (y-1) but on the other side you didn't distribute the 4 in to the x?

OpenStudy (anonymous):

I still can't figure this out =(

OpenStudy (anonymous):

@khuangg is wrong.

OpenStudy (anonymous):

@miss_karina Have you learnt the topic Indices?